OpenStudy (anonymous):

help again What is the area of a regular hexagon with side length of 24cm? pleasse

OpenStudy (tkhunny):

Chop it up. 1) Label one side 24 cm 2) Draw the two radii intersecting that side. 3) Label the two EQUAL base angles of the Isosceles triangle just drawn. Any ideas, yet?

OpenStudy (anonymous):

area of a regular hexagon=3rt3/2 *a^2 =3rt3/2*24^2 =72*12rt3 =864rt3 cm^2